The region R is bounded by the curve "y =4-x\u00b2" and the line "y=4-2x"


   a. Find the intersection points.

b.Sketch and shade the region R.


1
Expert's answer
2021-05-11T09:24:53-0400

a. intersection points we can find from equation


"4-x^2=4-2x"


"x^2-2x=0"


"x(x-2)=0"


"x_1=0"


"x_2=2"


"y_1=4-x_1^2=4-0=4"


"y_2=4-x_2^2=4-4=0"


Answer: intersection points: (0, 4), (2, 0).


b.